# arXiv:2603.01575v1[quant-ph] 02 Mar 2026

### 02 Mar 2026 v1

@@ Abstract
From the modern perspective, the measurement of a physical observable is regarded as a special case of POVMs, namely, Projection-Valued Measures (PVMs). However, PVMs are defined in an algebraic way, and their operational characterization remains unclear. The aim of this Letter is to provide an operational characterization of PVMs among POVMs. In order to verify a genuinely operational interpretation of the results, we employ the framework of Generalized Probabilistic Theories (GPTs) [1, 2, 3, 4, 5, 6, 7, 8, 9], which retain only the operational requirements regarding states and measurements from the axioms of quantum theory while discarding the rest. Within GPTs, the generalization of projective measurements, and therefore also that of “physical observables,” is not yet clear, and this question remains an active topic of research [10, 11, 12, 13, 14, 15, 16, 17]. Numerous studies have also attempted to quantify the “PVM-likeness” of a POVM [18, 19, 20, 21]; however, these approaches often lack a clear operational interpretation or fail to provide a complete characterization of PVMs.
In this Letter, we provide a fully operational characterization of PVMs along with their quantification by generalizing the concept of intersubjectivity introduced by Ozawa [22] to GPTs. A quantum measurement is called intersubjective if two observers performing the measurement simultaneously are guaranteed to obtain the same outcome. Ozawa argued that such inter-observer agreement is essential for a measurement to ascertain an observer-independent value, rather than merely generating outcomes randomly (Fig. 1). While Ozawa proved that every PVM is intersubjective in quantum theory [22], the converse implication remains an open question.
@@
A finite-outcome measurement is given by a family $A=(a_{x})_{x\in X}$ of effects satisfying $\sum_{x\in X}a_{x}=1_{S}$. When a measurement $A=(a_{x})_{x\in X}$ is performed on a state $s\in S$, the probability of obtaining the outcome $x\in X$ is $a_{x}(s)$. We denote by $\mathcal{M}_{S}(X)$ the set of all measurements with outcome set $X$. We extend several of the results for continuous-outcome measurements; see Section S3 of the Supplemental Material.
@@
We now show that the conditions of intersubjectivity and sharpness, although conceptually distinct at first glance, are, in fact, equivalent in any physical theory. Moreover, they are quantitatively related to each other.
Theorem 1.
  • (i)
    If an $n$-outcome measurement is $\alpha$-sharp, then it is $\bigl(1-\bigl(n^{2}-n\bigr)(1-\alpha)\bigr)$-intersubjective.
  • (ii)
    If a measurement is $\alpha$-intersubjective, then it is $\frac{1+\alpha}{2}$-sharp.
In particular, a measurement is intersubjective if and only if it is sharp.
The proof is given in the End Matter. In the case of two-outcome measurements, (i.e., $n=2$), the two quantifications are equivalent: a two-outcome measurement is $\alpha$-intersubjective if and only if it is $\frac{1+\alpha}{2}$-sharp.
@@
Although Ozawa [22] proved that all PVMs are intersubjective and that not all POVMs are intersubjective, it remained an open question whether there exists an intersubjective POVM other than PVMs. By applying Theorem 1 to quantum theory, we find the following theorem.
@@ A POVM $A=(a_{x})_{x\in X}$ is intersubjective if and only if, for any $x\neq x^{\prime}\in X$,
Next, we discuss the basic properties of intersubjectivity. First, we describe its behavior under joint measurements and probabilistic mixtures.
Proposition 4.
@@
    Let $C=\lambda A+(1-\lambda)B$ with $\lambda\in(0,1]$ be a probabilistic mixture of two measurements $A=(a_{x})_{x\in X}$ and $B=(b_{x})_{x\in X}$. If $C$ is (completely) $\alpha$-intersubjective, then $A$ is (completely) $\bigl(1-\frac{1-\alpha}{\lambda}\bigr)$-intersubjective.
Here, the probabilistic mixture of two measurements $A=(a_{x})_{x\in X}$ and $B=(b_{x})_{x\in X}$ is defined as $\lambda A+(1-\lambda)B:=(\lambda a_{x}+(1-\lambda)b_{x})_{x\in X}$ for $\lambda\in[0,1]$. The proof is given in the End Matter.
As a special case of Proposition 4-(i), a joint measurement of intersubjective measurements must be intersubjective. However, a joint measurement of two completely intersubjective measurements is not always completely intersubjective in general, even though this holds in classical and quantum theories [50, 51, 52]. An explicit counterexample is presented in Example 4 in the End Matter.
Proposition 4-(ii) implies that adding noise does not improve the intersubjectivity of a measurement. This suggests that the quantification of (complete) intersubjectivity can also be interpreted as quantitative indicators of “noiselessness” of measurements. Importantly, this interpretation aligns with the historical perspective: PVMs have traditionally been regarded as perfectly accurate, while POVMs are generalized measurements that may incorporate noise. Accordingly, the quantification of “PVM-likeness” has also been discussed in the context of measuring noise [18, 19, 20, 21].
Proposition 4 demonstrates the importance of the degree of (complete) intersubjectivity. However, determining this value for a given measurement is generally nontrivial. Here, we present results for several simple cases. A coin-tossing measurement $A=(\lambda_{x}1_{S})_{x\in X}$ is (completely) $\alpha$-intersubjective if and only if
@@
The proofs of these results are provided in the Supplemental Material.
@@
State discrimination is the task of distinguishing a family of states through a single-shot measurement: given a finite ensemble of states $(s_{x},p_{x})_{x\in X}$, where state $s_{x}$ is prepared with probability $p_{x}$, the goal is to identify which state has been prepared. One performs a measurement $A=(a_{x})_{x\in X}$ on the given state and guesses that the prepared state was $s_{x}$ when outcome $x$ is obtained. The success probability of this task is given by $\sum_{x\in X}p_{x}a_{x}(s_{x})$. A family of states $(s_{x})_{x\in X}$ is said to be perfectly distinguishable by a measurement $A=(a_{x})_{x\in X}$ if $a_{x}(s_{x})=1$ holds for all $x\in X$.
We observe the following.
Theorem 6.
@@
    Every completely intersubjective measurement admits a family of states that is perfectly distinguishable by it.
Statement (i) follows from the fact that every extremal two-outcome measurement is sharp [12], while (ii) follows from Bauer’s maximum principle and the implication extremal $\implies$ intersubjective. Statement (iii) follows from the implication completely intersubjective $\implies$ elementwise sharp, along with the fact that for any sharp effect $a\neq 0_{S}$, there exists a state $s$ such that $a(s)=1$ [12].
@@
The authors thank the Yukawa Institute for Theoretical Physics at Kyoto University, where this work was initiated during the YITP-W-25-11. We thank Hiroatsu Okamitsu, Masanao Ozawa, Ryo Takakura, and Kaito Watanabe for fruitful discussions. S. U. was supported by Forefront Physics and Mathematics Program to Drive Transformation, a World-leading Innovative Graduate Study Program, the University of Tokyo. K. O. was supported by Advanced Basic Science Course, a World-leading Innovative Graduate Study Program, the University of Tokyo. H. A. was supported by JSPS KAKENHI Grant Number 25KJ0043.
@@
By Theorem 1, it suffices to show that for any positive operators $a,b\geq 0$, the following two conditions are equivalent:
@@
Appendix S2 Detailed calculations of intersubjectivity
In this appendix, we provide proofs of the values of (complete) intersubjectivity in several simple cases.
@@
Several of our results presented in the main text of the Letter (Theorem 3, Proposition 4 without quantifications, and the implication extremal $\implies$ intersubjective) remain valid for continuous-outcome measurements. In this appendix, we extend our analysis to this more general setting.
@@
Proposition S.7 (Continuous-outcome version of Proposition 4-(i)).
@@
Proposition S.8 (Continuous-outcome version of Proposition 4-(ii)).

### 28 May 2026 v2

# Intersubjectivity as a principle determining physical observables and non-classicality

@@ Abstract
From the modern perspective, the measurement of a physical observable is regarded as a special case of POVMs, namely, Projection-Valued Measures (PVMs). However, PVMs are defined in an algebraic way, and their operational characterization remains unclear. The aim of this Letter is to provide an operational characterization of PVMs among POVMs. In order to verify a genuinely operational interpretation of the results, we employ the framework of Generalized Probabilistic Theories (GPTs) [1, 2, 3, 4, 5, 6, 7, 8, 9], which retain only the operational requirements regarding states and measurements from the axioms of quantum theory while discarding the rest. Within GPTs, the generalization of projective measurements, and therefore also that of “physical observables,” is not yet clear, and this question remains an active topic of research [10, 11, 12, 13, 14, 15, 16, 17]. Numerous studies have also attempted to quantify the “PVM-likeness” of a POVM [18, 19, 20, 21, 22]; however, these approaches often lack a clear operational interpretation or fail to provide a complete characterization of PVMs.
In this Letter, we provide a fully operational characterization of PVMs along with their quantification by generalizing the concept of intersubjectivity [23, 24, 25] introduced by Ozawa to GPTs. A quantum measurement is called intersubjective if two observers performing the measurement simultaneously are guaranteed to obtain the same outcome. Ozawa argued that such inter-observer agreement is essential for a measurement to ascertain an observer-independent value, rather than merely generating outcomes randomly (Fig. 1). While Ozawa proved that every PVM is intersubjective in quantum theory [23], the converse implication remains an open question.
@@
A finite-outcome measurement is given by a family $A=(a_{x})_{x\in X}$ of effects satisfying $\sum_{x\in X}a_{x}=1_{S}$. When a measurement $A=(a_{x})_{x\in X}$ is performed on a state $s\in S$, the probability of obtaining the outcome $x\in X$ is $a_{x}(s)$. We denote by $\mathcal{M}_{S}(X)$ the set of all measurements with outcome set $X$. We extend several of the results for continuous-outcome measurements; see the Supplemental Material  [^1] See Supplemental Material for further details. .
@@
We now show that the conditions of intersubjectivity and sharpness, although conceptually distinct at first glance, are, in fact, equivalent in any physical theory:
Proposition 1.
A measurement is intersubjective if and only if it is sharp.
Moreover, this equivalence admits a quantitative refinement, which is presented in the End Matter.
@@
Although Ozawa [23] proved that all PVMs are intersubjective and that not all POVMs are intersubjective, it remained an open question whether there exists an intersubjective POVM other than PVMs. By applying Proposition 1 to quantum theory, we find the following theorem.
@@ A POVM $A=(a_{x})_{x\in X}$ is intersubjective if and only if, for any $x\neq x^{\prime}\in X$,
Next, we discuss the basic properties of the degree of intersubjectivity. First, we describe its behavior under joint measurements and probabilistic mixtures:
@@
    If measurements $A$ and $B$ are $\alpha$- and $\beta$-intersubjective, respectively, then any joint measurement $C\in\mathrm{JM}(A,B)$ is $(\alpha+\beta-1)$-intersubjective.
@@
    Consider a probabilistic mixture $C=\lambda A+(1-\lambda)B$ with $\lambda\in(0,1]$, where $A=(a_{x})_{x\in X}$ and $B=(b_{x})_{x\in X}$. If $A$ is not (completely) $\alpha$-intersubjective, then $C$ is not (completely) $(1-\lambda(1-\alpha))$-intersubjective.
Here, the probabilistic mixture of two measurements $A=(a_{x})_{x\in X}$ and $B=(b_{x})_{x\in X}$ is defined by $\lambda A+(1-\lambda)B:=(\lambda a_{x}+(1-\lambda)b_{x})_{x\in X}$ for $\lambda\in[0,1]$. The proof is given in the Supplemental Material footnotemark: .
As a special case of statement (i), a joint measurement of intersubjective measurements must be intersubjective. However, a joint measurement of two completely intersubjective measurements is not always completely intersubjective in general, even though this holds in classical and quantum theories [54, 55, 56]. An explicit counterexample is presented in Example 4 in the End Matter.
Statement (ii) implies that adding noise does not improve the intersubjectivity of a measurement. This suggests that the quantification of (complete) intersubjectivity can also be interpreted as quantitative indicators of “noiselessness” of measurements. Importantly, this interpretation aligns with the historical perspective: PVMs have traditionally been regarded as perfectly accurate, while POVMs are generalized measurements that may incorporate noise. Accordingly, the quantification of “PVM-likeness” has also been discussed in the context of measuring noise [18, 19, 20, 21, 22].
These observations demonstrate the importance of the degree of (complete) intersubjectivity. However, evaluating this quantity for a given measurement is generally nontrivial. We now present results for several simple cases. A coin-tossing measurement $A=(\lambda_{x}1_{S})_{x\in X}$ is (completely) $\alpha$-intersubjective if and only if
@@
The proofs of these results are provided in the Supplemental Material footnotemark: .
@@
State discrimination is the task of distinguishing a family of states through a single-shot measurement: given a finite ensemble of states $(s_{x},p_{x})_{x\in X}$, where state $s_{x}$ is prepared with probability $p_{x}$, the goal is to identify which state has been prepared. One performs a measurement $A=(a_{x})_{x\in X}$ on the given state and guesses that the prepared state was $s_{x}$ when outcome $x$ is obtained. The success probability of this task is given by $\sum_{x\in X}p_{x}a_{x}(s_{x})$. A family of states $(s_{x})_{x\in X}$ is said to be perfectly distinguished by a measurement $A=(a_{x})_{x\in X}$ if $a_{x}(s_{x})=1$ holds for all $x\in X$.
We observe the following:
@@
    Every completely intersubjective measurement admits a family of states that is perfectly distinguished by it.
Statement (i) follows from the fact that every extremal two-outcome measurement is sharp [12], while (ii) follows from Bauer’s maximum principle and the implication extremal $\implies$ intersubjective. Statement (iii) follows from the implication completely intersubjective $\implies$ elementwise sharp, along with the fact that for any sharp effect $a\neq 0_{S}$, there exists a state $s$ such that $a(s)=1$ [12]. We note, however, that unlike in the quantum case, a perfectly distinguishable family of states need not be perfectly distinguished by a completely intersubjective measurement, as demonstrated by Example 4 in the End Matter.
@@
The authors thank the Yukawa Institute for Theoretical Physics at Kyoto University, where this work was initiated during the YITP-W-25-11. We thank Hiroatsu Okamitsu, Masanao Ozawa, Ryo Takakura, Kaito Watanabe, and Ludovico Lami for fruitful discussions. S. U. was supported by Forefront Physics and Mathematics Program to Drive Transformation, a World-leading Innovative Graduate Study Program, the University of Tokyo. K. O. was supported by Advanced Basic Science Course, a World-leading Innovative Graduate Study Program, the University of Tokyo. H. A. was supported by JSPS KAKENHI Grant Number 25KJ0043 and 26K17031.
@@
By Proposition 1, it suffices to show that for any positive operators $a,b\geq 0$, the following two conditions are equivalent:
@@
Appendix S2 Proofs of quantitative properties of intersubjectivity
In this appendix, we prove the quantitative properties of intersubjectivity stated in the main text. We begin with its behavior under joint measurements and probabilistic mixtures.
Proposition S.2.
  • (i)
  • (ii)
    Let $C=\lambda A+(1-\lambda)B$ with $\lambda\in(0,1]$ be a probabilistic mixture of two measurements $A=(a_{x})_{x\in X}$ and $B=(b_{x})_{x\in X}$. If $A$ is not (completely) $\alpha$-intersubjective, then $C$ is not (completely) $(1-\lambda(1-\alpha))$-intersubjective.
Proof.
  • (i)
    Let $D=(d_{x,y,x^{\prime},y^{\prime}})_{(x,y),(x^{\prime},y^{\prime})\in X\times Y}\in\mathrm{JM}(C,C)$. Define two measurements $E=(e_{x,x^{\prime}})_{x,x^{\prime}\in X}$ and $F=(f_{y,y^{\prime}})_{y,y^{\prime}\in Y}$ by
    $e_{x,x^{\prime}}:=\sum_{y,y^{\prime}}d_{x,y,x^{\prime},y^{\prime}},\quad f_{y,y^{\prime}}:=\sum_{x,x^{\prime}}d_{x,y,x^{\prime},y^{\prime}}.$
    $\sum_{x\in X}e_{x,x}\geq\alpha 1_{S},\quad\sum_{y\in Y}f_{y,y}\geq\beta 1_{S}.$
    Therefore,
    $\sum_{(x,y)\in X\times Y}d_{x,y,x,y}$ $=\sum_{x,y,y^{\prime}}d_{x,y,x,y^{\prime}}+\sum_{x,y,x^{\prime}}d_{x,y,x^{\prime},y}$
    $\qquad{}-\sum_{\begin{subarray}{c}x,y,x^{\prime},y^{\prime}\\ x=x^{\prime}\text{ or }y=y^{\prime}\end{subarray}}d_{x,y,x^{\prime},y^{\prime}}$
    $\geq(\alpha+\beta-1)1_{S}.$
    This completes the proof.
  • (ii)
    We first prove the statement for intersubjectivity. Suppose that $C$ is $(1-\lambda(1-\alpha))$-intersubjective. Let $D=(d_{x,x^{\prime}})_{x,x^{\prime}\in X}\in\mathrm{JM}(A,A)$ and $E=(e_{x,x^{\prime}})_{x,x^{\prime}\in X}\in\mathrm{JM}(B,B)$. Then we have $\lambda D+(1-\lambda)E\in\mathrm{JM}(C,C)$. Thus,
    $\sum_{x\in X}\bigl(\lambda d_{x,x}+(1-\lambda)e_{x,x}\bigr)\geq(1-\lambda(1-\alpha))1_{S}.$
    Therefore,
    $\sum_{x\in X}d_{x,x}\geq\frac{1-\lambda(1-\alpha)}{\lambda}1_{S}-\frac{1-\lambda}{\lambda}\sum_{x\in X}e_{x,x}\geq\alpha 1_{S}.$
Next, we present detailed calculations of the degree of (complete) intersubjectivity in several simple cases.
@@
Appendix S3 Proofs of Example 3 and Example 4
In this appendix, we provide proofs of Example 3 and Example 4 presented in the End Matter.
Proof of Example 3.
Let
$S:=\left\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}\>\middle|\>-x_{3}\leq x_{1}\leq x_{3},\ -x_{3}\leq x_{2}\leq x_{3},\ x_{3}=1\right\},$
and let $A:=\frac{1}{2}B+\frac{1}{2}C$, where $B=(b_{+},b_{-})$ and $C=(c_{+},c_{-})$ are defined by
$b_{\pm}(x_{1},x_{2},x_{3}):=\frac{1}{2}(x_{3}\pm x_{1}),\quad c_{\pm}(x_{1},x_{2},x_{3}):=\frac{1}{2}(x_{3}\pm x_{2}).$
We show that $A$ is completely intersubjective but not extremal. It is clear that $A$ is not extremal. Hence, it suffices to show that $A$ is sharp. Suppose that an effect $d$ satisfies
$d\leq\frac{1}{2}b_{+}+\frac{1}{2}c_{+},\quad d\leq\frac{1}{2}b_{-}+\frac{1}{2}c_{-}.$
Evaluating these inequalities at the states (1,1,1)(-1,-1,1) and $(1,1,1)$, respectively, gives d(1,1,1)=d(1,1,1)=0d(-1,-1,1)=d(1,1,1)=0. Moreover, since $d$ is affine, we have d(1,1,1)+d(1,1,1)=d(1,1,1)+d(1,1,1)=0d(1,-1,1)+d(-1,1,1)=d(-1,-1,1)+d(1,1,1)=0. By positivity of $d$, this implies d(1,1,1)=d(1,1,1)=0d(1,-1,1)=d(-1,1,1)=0. Therefore, $d=0_{S}$, and hence $A$ is sharp. ∎
Proof of Example 4.
Let
$S:=\left\{(x_{1},x_{2},x_{3},x_{4},x_{5})\in\mathbb{R}^{5}\>\middle|\>0\leq x_{i}\ (i=1,\dots,5),\ x_{5}\leq x_{1}+x_{4},\ x_{5}\leq x_{2}+x_{3},\ x_{1}+x_{2}+x_{3}+x_{4}=1\right\}.$
Let $A=(a_{1},a_{2},a_{3},a_{4})$, where
$a_{i}(x_{1},x_{2},x_{3},x_{4},x_{5}):=x_{i}\quad(i=1,2,3,4).$
We also define $B:=(a_{1}+a_{2},a_{3}+a_{4})$, $C:=(a_{1}+a_{3},a_{2}+a_{4})$, and $s_{i}:=(\delta_{1,i},\delta_{2,i},\delta_{3,i},\delta_{4,i},0)\in S$ for $i=1,2,3,4$.
We first show that $A$ is elementwise sharp and extremal, but not completely intersubjective. To prove elementwise sharpness, suppose that an effect $e$ satisfies $e\leq a_{1}$ and $e\leq 1_{S}-a_{1}$. Then $e$ must be of the form
$e(x_{1},x_{2},x_{3},x_{4},x_{5})=\lambda x_{5}$
for some constant $\lambda$. Positivity of $e$ gives $\lambda\geq 0$, while evaluating $e\leq a_{1}$ at $\bigl(0,0,\frac{1}{2},\frac{1}{2},\frac{1}{2}\bigr)\in S$ gives $\lambda\leq 0$. Thus $\lambda=0$, and hence $e=0_{S}$. Therefore $a_{1}$ is a sharp effect. The same argument applies to $a_{2}$, $a_{3}$, and $a_{4}$.
The extremality of $A$ follows from the fact that $a_{1},a_{2},a_{3},a_{4}$ are linearly independent indecomposable effects.
On the other hand, $A$ is not completely intersubjective. Indeed, the effect $e$, defined by
$e(x_{1},x_{2},x_{3},x_{4},x_{5}):=x_{5},$
satisfies $e\leq a_{1}+a_{4}$ and $e\leq a_{2}+a_{3}$. Thus the coarse-graining $D:=(a_{1}+a_{4},a_{2}+a_{3})$ of $A$ is not sharp, and hence $A$ is not completely intersubjective.
We next show that $B$ and $C$ are completely intersubjective. Suppose that an effect $e$ satisfies $e\leq a_{1}+a_{2}$ and $e\leq a_{3}+a_{4}$. Then $e$ must be of the form
$e(x_{1},x_{2},x_{3},x_{4},x_{5})=\lambda x_{5}$
for some constant $\lambda$. Positivity of $e$ gives $\lambda\geq 0$, while evaluating $e\leq a_{1}+a_{2}$ at $\bigl(0,0,\frac{1}{2},\frac{1}{2},\frac{1}{2}\bigr)\in S$ gives $\lambda\leq 0$. Thus $\lambda=0$, and hence $e=0_{S}$. Therefore, $B$ is sharp, and hence completely intersubjective. The same argument applies to $C$.
Finally, we show that $(s_{1},s_{2},s_{3},s_{4})$ is perfectly distinguished only by $A$. Let $A^{\prime}=(a^{\prime}_{1},a^{\prime}_{2},a^{\prime}_{3},a^{\prime}_{4})$ be a measurement that perfectly distinguishes these states. Then each $a^{\prime}_{i}$ must be of the form
$a^{\prime}_{i}(x_{1},x_{2},x_{3},x_{4},x_{5})=x_{i}+\lambda_{i}x_{5}$
for some constant $\lambda_{i}$. Positivity of $a^{\prime}_{i}$ implies $\lambda_{i}\geq 0$ for each $i$. Since $A^{\prime}$ is a measurement, we have $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=0$. Therefore $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=0$, which implies $A^{\prime}=A$. ∎
@@
Several of our results presented in the main text remain valid for continuous-outcome measurements. In this appendix, we extend our analysis to this more general setting.
@@
Proposition S.8.
@@
Proposition S.9.